OpenStudy (anonymous):

definite integral, lnx/sqrt((x-1)^3)

OpenStudy (anonymous):

\[\int\limits_{1}^{2}\ln(x)/\sqrt((x-1)^3)\]

OpenStudy (freckles):

Have you tried integration by parts? Not saying it will give us what we want but that is what I see worth trying right now.

OpenStudy (freckles):

Seems to work and then after that a sub might be required.